OpenStudy (anonymous):

The first term of an arithmetic sequence is 6, the third is 17. Find the 30th.

OpenStudy (anonymous):

do you know the equation for this?

OpenStudy (anonymous):

yeah isnt it an=a+(n-1)d

OpenStudy (anonymous):

here a=6 & a+2d=17,therefore u have 'd',nd then u hav 2 find a+29d..

OpenStudy (anonymous):

where do we get the 2 from?

OpenStudy (anonymous):

Adding on to what @campbell_st said: Since our formula is: \[ a_n = a_1+(n-1)d \]And we are given \( a_1=6,\;a_3=17\) we can make a system of equation(s): \[ a_3= 17 =6 + (3-1)d \]Then you solve for \(d\)

OpenStudy (anonymous):

then find a+29d for n=30..

OpenStudy (anonymous):

would it be 6+2d?

OpenStudy (anonymous):

cuz thats the 3rd term right?

OpenStudy (anonymous):

i got 11/2

OpenStudy (anonymous):

why is Tn? not really

OpenStudy (anonymous):

oh so its the same thing as an?

OpenStudy (anonymous):

okay so then i plug it in?

OpenStudy (anonymous):

would it be an=6+11/2?

OpenStudy (anonymous):

okay that helped a bit, so would it be an=15.31?

OpenStudy (anonymous):

but what if i cant use the calculator

OpenStudy (anonymous):

but what if i cant use any type of calculator tho/

OpenStudy (anonymous):

to the n-1?

OpenStudy (anonymous):

it was 30-1 and i got 29 which i timed it by 5.5 which i got 15.95

OpenStudy (anonymous):

but i thought you said 5.5 . where did 5 and .05 come from?

OpenStudy (anonymous):

yeah , where it says you multiply 29 times 5.5 5 and 0.5

OpenStudy (anonymous):

?

OpenStudy (anonymous):

29 *5.5 can't be 15.95. If you just drop the .5 from 5.5 and multiply by 29 you are going to get way more than 15.

OpenStudy (anonymous):

okay, but why is it that you drop the 5 from 5.5?

OpenStudy (anonymous):

To make your computation easier. We will multiply by .5 afterwards and add the result to what you get when you multiply 29 by 5.

OpenStudy (anonymous):

29*5+29*0.5

OpenStudy (anonymous):

is it cause it equals the same thing?

OpenStudy (anonymous):

except one has a decimal and one doesn't/

OpenStudy (anonymous):

when multiplying 29 and 5 you get what?

OpenStudy (anonymous):

you get 145?

OpenStudy (anonymous):

ok...good...now when you multiply 29 by 0.5 you get what?

OpenStudy (anonymous):

14.5?

OpenStudy (anonymous):

ok...so now add those two results together to get your final answer.

OpenStudy (anonymous):

would it be 29?

OpenStudy (anonymous):

How are you getting 29 when 29*5 is 145?

OpenStudy (anonymous):

i mean 39

OpenStudy (anonymous):

wait nevermind

OpenStudy (anonymous):

Do you bnot remember the rule for adding two numbers with decimals?

OpenStudy (anonymous):

not*

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

would it be 159.5?

OpenStudy (anonymous):

So, so far, you havee given one answer of 15.95, and now you have given me a 29 or a 39.

OpenStudy (anonymous):

Finally!

OpenStudy (anonymous):

There you go.

OpenStudy (anonymous):

okay but can you explain to me why there are 5 and .5 tho?

OpenStudy (anonymous):

Can't be too dependent on that calculator. Never know when you are going to be in a class where you can't use it.

OpenStudy (anonymous):

 Just another way to show that 29*5.5 =159.5 by splitting it up into 29*5 +29*.5